OpenStudy (anonymous):

Patrick counted the cars that drove by his house over a 5-minute period. He counted a total of 27 cars. Suppose that at the same time of the day exactly one week later, patrick counts cars over a 20-minute period. About how many cars would you expect him to count?

OpenStudy (anonymous):

about 20/5 * 27 = 4 * 27 = 108